If you had 12.00 grams of water at 70.0C, under constant pressure of 1.00 atm, how much energy would be released to cool it to -20.5C


1
Expert's answer
2021-03-29T06:03:23-0400

q = m c "\Delta T"

q = heat released

c = specific heat of water = 4.186 j/g⁰C

"\Delta T" = T2 - T1 = -20.5⁰C - 70⁰C = -95⁰C

m = 12 g

q = 12×4.186×(-95)

q = -4772.04 joules

q = 4772.04 joules released